jennylove:

2) In a closed system, Glider A with a mass of 0.50 kg and a speed of 1.50 m/s collides with Glider B at rest with a mass of 0.30 kg. The two interlock and move off. What speed are they moving? 1.3125 = 1.31 m/sec

jennylove:

i think its 1.31 ^

ramen:

i think so?

ramen:

idk yet

kittybasil:

Closed system = conservation of momentum. Equation:\[\Large{m_1u_1+m_2u_2=m_1v_1+m_2v_2}\]When the body is at rest, \(u_2=0\). Also, \(v_1=v_2=v\) since both interlock and move together. As a result, your equation is now\[\large{m_1u_1+m_2(0)=m_1v+m_2v}\]\[m_1u_1+0=m_1v+m_2v\]We know that your values are as follows:\[m_1=0.50\text{kg, }m_2=0.30\text{ kg, }u_1=1.50\text{ m/s}\]

kittybasil:

Plugging in those values, we get\[(0.50\text{ kg})(1.50\text{ m/s})+0=(0.50\text{ kg})v+(0.30\text{ kg})v\]Simplifying, we now have\[0.50(1.50)=0.50v+0.30v\]We can reverse distribute (source: Distributive Property) the right side to isolate the variable \(v\):\[0.50\cdot1.50=v(0.50+0.30)\]Simplifying again,\[0.75=v(0.80)\]As a result, your variable should be the result of this.\[0.75\div0.80=v\]

kittybasil:

Edit: sorry, velocity variable \(v\) should be your "speed" that the question is asked for. If you solve that expression it should give you the answer

kittybasil:

@jennylove Everything alright? If you have any questions, feel free to ask for further assistance :)

jennylove:

@kittybasil wow thank you for this deep explanation ! I found the answer to be 0.9375 m/s by using the info you gave me , am i correct?

kittybasil:

Yeah, that should be correct. But if you need to round up or down, don't forget to do that! :)

jennylove:

Okay, thank you so much! youve been very helpful